A student studying the weather for d days observed that (i) it rained on 7 days, morning or afternoon; (ii) When it rained in the afternoon, it was clear in the morning; (iii) there were five clear afternoons; and (iv) there were six clear mornings. then d equals :
(A) 7
(B) 11
(C) 10
(D) 9

Grade:11

1 Answers

Arun
5 years ago
d--1-2-3-4-5-6-7-8-9-10-11
am-c-r-c-r-c-r-c-c-c------
pm-r-c-r-c-r-c-r-c-c------
 
((5+6)+7)/2=18/2=9
add partial clear 5+6=11 and 7 rainy days for a total of 18 ms and pms
divide 18/2 to get 9 days